OpenStudy (anonymous):

Prove secx-1/tanx= tanx/secx+1

hero (hero):

Begin with the LHS: \[\frac{\sec(x) - 1}{\tan(x)}\] Multiply by \(\frac{\sec(x) + 1}{\sec(x) + 1}\): \[\frac{\sec(x) - 1}{\tan(x)} \dot\ \frac{\sec(x) + 1}{\sec(x) + 1}\] \[\frac{(\sec(x) - 1)(\sec(x) + 1)}{\tan(x)(\sec(x) + 1)}\] Notice that \((\sec(x) - 1)(\sec(x) + 1) = \sec^2x - 1\) since \((a + b)(a - b) = a^2 - b^2\) \[\frac{\sec^2x - 1}{\tan(x)(\sec(x) + 1)}\] Remember that \(\sec^2x - 1 = \tan^2x\) So \[\frac{\tan^2x}{\tan(x)(\sec(x) + 1)}\] Now simplify